# supplementary

### v3

@@ Abstract
Scaling up deliberative and voting participation is a longstanding endeavor – a cornerstone for direct democracy and legitimate collective choice. Recent breakthroughs in generative artificial intelligence (AI) and large language models (LLMs) unravel new capabilities for AI personal assistants to overcome cognitive bandwidth limitations of humans, providing decision support or even direct representation of human voters at large scale. However, the quality of this representation and what underlying biases manifest when delegating collective decision making to LLMs is an alarming and timely challenge to tackle. By rigorously emulating with high realism more than >50K LLM voting personas in 81 real-world voting elections, we disentangle the nature of different biases in LLMS (GPT 3, GPT 3.5, and Llama2). Complex preferential ballot formats exhibit significant inconsistencies compared to simpler majoritarian elections that show higher consistency. Strikingly though, by demonstrating for the first time in real-world a proportional representation of voters in direct democracy, we are also able to show that fair ballot aggregation methods, such as equal shares, prove to be a win-win: fairer voting outcomes for humans with fairer AI representation. This novel underlying relationship proves paramount for democratic resilience in progressives scenarios with low voters turnout and voter fatigue supported by AI representatives: abstained voters are mitigated by recovering highly representative voting outcomes that are fairer. These interdisciplinary insights provide remarkable foundations for science, policymakers and citizens to develop safeguards and resilience for AI risks in democratic innovations.
@@ 1 Introduction
Recent advances in artificial intelligence (AI) provide new unprecedented opportunities for citizens to scale up participation in digital democracy [1, 2, 3]. Generative AI in particular, such as large language models (LLMs), have a potential to overcome human cognitive bandwidth limitations and digitally assist citizens to deliberate and decide about public matters at scale [4, 5, 6, 7, 8]. This is by articulating, summarizing and even providing syntheses of complex opinions [9, 10, 7, 8], with a potential to mitigate for the voter fatigue and reduced voters turnout [11, 10], while fostering common ground for compromises, consensus and lower polarization [11, 12, 10, 13, 4]. However, understanding the implications and risks of using large language models for decision support/recommendations or even direct representation of human voters is a pressing challenge [14, 15, 16, 17]. In this article we study the inconsistencies of large language models to approximate both individual and collective choices as well as the inconsistencies of large language models across different preference elicitation methods. We explore causal links of these inconsistencies to potential cognitive biases triggered by the input to large language models based on which choices are made.
Generative AI voting: a converging advance with inevitable challenges. Large language models have been used to predict election winners using sensitive demographic data about the political profile of individuals [18]. They have also been used to predict pairwise comparisons of proposals for constitutional changes [6] and summarized opinions expressed with free-form text [19]. However, little is known whether this AI predictive capability can expand to more complex collective decision-making processes that involve more options to choose from, with more complex preference elicitation methods [5].
Participatory budgeting [20] is one such process put under scrutiny in this article. Here city authorities distribute a public budget by letting citizens propose their own project ideas, which they vote for and often implement themselves [21]. Projects may be pertinent to different impact areas (e.g. environment, culture, welfare), beneficiaries (e.g. elderly, children) and can have different costs. Voters can approve, rank or distribute points over their preferred projects, while winners are elected based on the popularity of the projects (utilitarian greedy) or based on a proportional representation of the voters’ preferences (equal shares[22, 23]. Recently though, a reinforcement learning approach has been used to elect winners by assisting voters to reach a consensus [24]. So far, the scope and citizens’ engagement in participatory budgeting campaigns remains to a large extent a one-shot and highly local [25, 11]. What if there was a way to geographically spread such decision-making processes and run them on a regular basis? With such complexity and degree of design freedom, scaling up participatory budgeting turns into the ultimate democratic blueprint to assess capabilities and risks of generative AI voting.
Inconsistencies among different preference elicitation methods, known also as intransitivity of preferences [26], have not been studied so far for large language models. However, there is evidence that such inconsistencies do manifest even in human voters [22, 23, 27], especially under polarized contexts [28] that are usually subject of biases [29, 30]. Whether potential biases that explain the inconsistencies between human and AI choices are of different nature than the ones between different input voting methods is an open question studied in this article.
How resilient representative voting outcomes are with generative AI. We hypothesize that generative AI representatives of human voters build up resilience for representative voting outcomes in voting processes of low turnout if these human voters would otherwise abstain or would not have capacity to actively participate (see Figure 1d). In other words, we explore whether the inconsistencies of voting outcomes because of low voting turnout are higher than inconsistencies originated from generative AI representatives of abstained voters.
The inconsistencies of generative AI, along with the potential biases explaining these inconsistencies, are systematically studied here for the first time using a novel factorial design based on real-world empirical evidence. It consists of 5 dimensions (see Figure 1): (i) Real-world voting scenarios - datasets from the 2012, 2016 and 2020 US national elections [31] as well as the mock and actual participatory budgeting voting of the ‘City Idea’ campaign of 2023 in Aarau, Switzerland [32] are studied. They come with different number of ballot options and voters, see Figure 1a and Section 4.2. (ii) Personal human traits - for each voter, multiple incremental levels of additional information are given as input to large language models to generate ballots. This information includes socio-demographics, political preferences, preferences for project categories and preferences for qualities of voting outcomes (popular vs. fair election of winners, individual vs. collective benefit). These are derived from real-world evidence including survey data collected along the voting campaigns (see Section 4.1, and Table LABEL:table:prologue1 and LABEL:table:prologue2). (iii) Voting input methods - four methods with incremental level of complexity and expressiveness are compared [28, 33]. These include single choice for all voting scenarios, approvals, score (assigning a preference score to each option) and cumulative voting (distributing a number of points over the options) [34, 35, 36] for the participatory budgeting scenarios. (iv) AI models - Three large language models are assessed along with a more mainstream predictive machine learning (ML) model used as a benchmark. GPT 3, GPT 3.5, and Llama2 are chosen covering a large spectrum of capabilities in open and closed generative AI [37], see Section 4.1 for more information. (v) Voting aggregation methods - majority is used for the scenario of US elections. The standard utilitarian greedy and the method of equal shares [38] are used for the participatory budgeting scenarios. The former method iterates to elect the next most popular project as long as the budget is not exhausted. It has been the standard method used in real-world. The latter method elects the projects that make the most people ‘happy’, meaning a combination of projects that maximizes a proportional representational of voters’ preferences. This is the method that was actually used in City Idea campaign to elect the winners [39], providing significance and realism for the findings of this study. See section 4.1 for more information.
The dimensions of the factorial design are illustrated in Table 1 and the studied combinations are marked with the colored boxes in Figure 1a. This broad spectrum of analysis based on real-world evidence allows us to generalize the findings of the study and make them highly relevant for a broad spectrum of research communities and policymakers.
@@
Figure 1: An overview of the generative AI framework, along with the factorial design and the biases studied in this article. (A) The factorial design with the 5 studied dimensions: (i) real-world voting scenarios in the context of participatory budgeting and national elections. (ii) Various combinations of persona human traits based on which AI personas are created. (iii) Four voting input methods. (iv) Four AI models, three large language models and a predictive machine learning model (benchmark). (v) Voting aggregation methods for elections and participatory budgeting. The studied combinations for each voting scenario are marked with different colors, see also Table 1. (B) The framework of generative AI voting. For each voter in the real-world voting scenario, a prompt is given to large language models to construct the voting persona. The input is the personal human traits, the voting options and the ballot format with instructions for the voting persona about how to make a choice. This choice is the output of the persona. Both human and AI choices are aggregated using a voting aggregation method. The inconsistencies of individual and collective choices for humans and AI personas are assessed, along with potential biases that explain these inconsistencies. (C) The personal human traits are mapped to cognitive biases. Section 4.4 illustrates choice inconsistencies and their origin to potential cognitive biases. (D) Hypothesizing for a recovery of consistency in voting outcomes under low voters turnout using AI representatives for abstained human voters.
@@
Assessing generative AI voting in action. Voting personas are constructed using input prompts of large language models as depicted in Figure 1b. This designed process aims to emulate the three voting scenarios with the different settings of Figure 1a. Each input prompt consists of a standardized description of the voter’s profile (see Section LABEL:sec:study in SI) and an instruction to vote according to the ballot format. The individual and collective real-world choices of humans and AI personas are compared for the first time by measuring the overlap (Jaccard similarity) of their selected/winning options as an indicator of consistency [28]. These consistency values are then becoming the dependent variable to predict using the personal human traits as independent variables (features), which are fed in a neural network (see Section 4.5). Based on a systematic mapping of human personal traits to the cognitive biases as illustrated in Figure 1c (see Section 4 for more detail), this prediction model causally explains the human traits that contribute to inconsistencies and the potential underlying biases that explain these inconsistencies. This novel analysis is designed to provide a significant conceptual advance to understand how voting design reinforces or mitigates different AI biases in real-world practice.
@@ 2 Results
1. Fair voting methods to elect winners are more resilient to human-AI inconsistencies, demonstrating a striking underlying win-win relationship: fairer voting outcomes for humans with fairer human representation by AI (Figure 2). These inconsistencies are particularly prominent in complex ballot formats with a large number of alternatives, while simple majoritarian voting tends to be highly consistent. GPT 3.5 shows transitivity of preferences among different ballot formats (Figure 3).
2. By representing half of the human voters who abstain with AI, a significant lost consistency is recovered in voting scenarios with low number of alternatives and fair voting aggregation methods to elect winners (Figure 4). For more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI in single-choice majoritarian voting with two alternatives.
3. Conformity biases explain consistency of human-AI choice, while unconscious biases explain inconsistency (Figure 5). Political/ideological profile determines the human-AI consistency in single-choice majoritarian voting with two candidates. Time discounting factors explain transitivity of AI preferences between different ballot formats, while unconscious biases explain intransitivity.
@@ 2.1 Fair collective choice is resilient to human-AI inconsistencies
Figure 2 illustrates the human-AI consistency in individual and collective choices. The consistency of individual choice remains poor in complex ballot formats with several alternatives. On average, it is 4.5% and 27.2% for the actual and mock voting in participatory budgeting, yet it is 82.3% for the binary majoritarian US elections. GPT 3.5 shows the highest consistency of individual choice among the large language models, which is 5.2% and 5.31% higher than GPT 3 and Llama2. On the contrary, the consistency of the collective choice increases by 51.2% in total. Strikingly, the consistency of equal shares is on average 80.8%, which is 32.8% higher than utilitarian greedy. Even when reducing the number winners of equal shares to the 8 of utilitarian greedy, the consistency remains 29.4% higher.
@@
Figure 2: Choice by large language models is consistent to humans for single-choice majoritarian elections, however consistency drops for more complex ballots with larger number of alternatives as in the case of participatory budgeting. Strikingly, consistency of collective choice is significantly higher than individual choice, particularly for the fairer voting aggregation rule of equal shares. GPT 3.5 shows the highest choice consistency and Llama2 the lowest among the large language models, which though remain inferior to a predictive machine learning model. The consistency (y-axis) in individual and collective choice is shown for different AI models (x-axis) across three real-world voting scenarios: The participatory budgeting campaign of City Idea, (A) actual and (B) mock, and (C) the US national elections of 2012, 2016 and 2020. The consistency in the participatory budgeting scenarios is shown for different voting input methods (left and right). The consistency of collective choice with the two voting aggregation methods of utilitarian greedy and equal shares is also shown. For the actual voting of City Idea, the consistency of equal shares is calculated for all winners and 8 winners (as many as greedy) for a fairer comparison.
Compared to large language models, the machine learning model shows 3.6% higher consistency in individual choice and 6.6% higher in collective choice. All consistency values are the maximum ones derived with all personal human traits as shown in Figure 1a. The removal of project preferences shows the highest consistency reduction of 19.7%, while political interest shows the lowest consistency reduction of 3.4%.
@@
Figure 3 illustrates the transitivity of preferences in different large language models by measuring the consistency of individual choice among different pairs of ballot formats. ‘Approval - cumulative’ voting shows the highest consistency, which is 1.8% higher than ‘single choice - approval’ and 3.5% higher than ‘single choice - cumulative’.
Figure 3: Voting personas based on GPT 3.5 show transitivity of the preferences over the different voting input methods in the actual complex voting scenario. However, GPT 3 and Llama2 remain with a low consistency <0.5. The simpler mock voting scenario shows higher transitivity among all language models. The consistency (y-axis) in individual choice among different pairs of ballot formats (x-axis) is shown for different large language models, humans and the two voting scenarios in the participatory budgeting campaign of City Idea: (A) actual vs. (B) mock.
GPT 3.5 shows 21.1% higher consistency than GPT 3 and 20.2% higher than Llama2. While human transitivity approaches 95.2% on average, AI transitivity remains at 70.1% for GPT 3.5. The mock voting with only 5 alternatives achieves on average 65.5% of consistency, while the actual voting with 33 alternatives shows a consistency of 43.4%.

### v4

@@ Abstract
Scaling up deliberative and voting participation is a longstanding endeavor – a cornerstone for direct democracy and legitimate collective choice. Recent breakthroughs in generative artificial intelligence (AI) and large language models (LLMs) unravel new capabilities for AI personal assistants to overcome cognitive bandwidth limitations of humans, providing decision support or even direct representation of abstained human voters at large scale. However, the quality of this representation and what underlying biases manifest when delegating collective decision making to LLMs is an alarming and timely challenge to tackle. By rigorously emulating with high realism more than >50K LLM voting personas in 306 real-world voting elections, we disentangle the nature of different biases in LLMs. Complex preferential ballot formats exhibit significant inconsistencies compared to simpler majoritarian elections that show higher consistency. Strikingly though, fair ballot aggregation methods, such as equal shares, prove to be a win-win: fairer voting outcomes for humans with fairer AI representation, especially for voters who are likely to abstain. This novel underlying relationship proves paramount for democratic resilience in progressive scenarios with low voters turnout by voter fatigue supported using AI representatives: abstained voters are mitigated by recovering highly representative voting outcomes that are fairer. These interdisciplinary insights provide remarkable foundations for science, policymakers and citizens to develop safeguards and resilience for AI risks in democratic innovations.
@@ 1 Introduction
Recent advances in artificial intelligence (AI) provide new unprecedented opportunities for citizens to scale up participation in digital democracy [1, 2, 3]. Generative AI in particular, such as large language models (LLMs), have a potential to overcome human cognitive bandwidth limitations and digitally assist citizens to deliberate and decide about public matters at scale [4, 5, 6, 7, 8]. This is by articulating, summarizing and even providing syntheses of complex opinions [9, 10, 7], with a potential to mitigate for the voter fatigue and reduced voters turnout [11, 10, 12], while fostering common ground for compromises, consensus and lower polarization [11, 13, 10, 14, 4]. However, understanding the implications and risks of using large language models for decision support, recommendations or even direct representation of human voters is a pressing challenge [15, 16, 17].
Unraveling inconsistencies in generative AI voting. We disentangle the inconsistencies of large language models in both individual and collective choices. In particular, we study three manifestations of choice inconsistency as shown in Figure 1a: (i) inconsistency by under-representation as a result of low human voters turnout, (ii) inconsistency by inaccurate approximation of human choice by AI and (iii) inconsistency by intransitivity [18] of AI choice over different ballot formats. Intransitivity even manifests among human voters, especially under polarized contexts [19] that are usually subject of biases [20, 21], therefore, we expect LLM choices to be vulnerable to such inconsistencies. Whether potential biases that explain the inconsistencies between human and AI choices are of different nature than the ones between different input voting methods is an open question studied in this article. We rigorously measure such inconsistencies with a single universal approach grounded to social choice theory [13, 22]. It exhaustively characterizes the similarity of the two choices (individual or collective) by counting the relative number of Condorcet pairwise matches, see Section 4.3 for further information. We also explore causal links of these inconsistencies to potential cognitive biases triggered by the input to large language models based on which choices are made.
@@
Figure 1: An overview of the studied generative AI voting framework. (A) Three manifestations of choice inconsistency are distinguished, measured using Condorcet pairwise matches: (i) inconsistencies by under-representation as a result of low voters turnout, (ii) inconsistencies by inaccuracy of AI choice to approximate human choice and (iii) inconsistency by intransitivity of AI choices over different ballot formats. (B) The factorial design with the 7 studied dimensions: (i) Real-world voting scenarios in the context of participatory budgeting and national elections. (ii) Various combinations of personal human traits (features) based on which AI voting personas are created. (iii) Four ballot formats. (iv) Four AI models, three large language models and a predictive machine learning model (benchmark). (v) Ballot aggregation methods for elections and participatory budgeting. (vi) The three abstaining models that are based for engagement, digital literacy and trust. (vii) Participation modalities ranging from exclusive human participation of varying turnout to mixed populations of humans and AI representatives of abstained voters. The studied combinations for each voting scenario are marked with different colors, see also Table 1. (C) The framework of generative AI voting. For each voter in the real-world voting scenario, a prompt is given to large language models to construct the voting persona. The input is the personal human traits, the voting options, and the ballot format, with instructions for the voting persona on how to make a choice. This choice is the output of the persona. Both human and AI choices are aggregated using a ballot aggregation method. The inconsistencies of individual and collective choices for humans and AI personas are assessed, along with potential biases that explain these inconsistencies. (D) The personal human traits are mapped to cognitive biases. Section LABEL:sec:bias illustrates the origin of choice inconsistencies to potential cognitive biases.
Generative AI voting: a converging advance technology with inevitable challenges. Large language models have been used to predict election winners using sensitive demographic data about the political profile of individuals [23]. They have also been used to predict pairwise comparisons of proposals for constitutional changes [6] and summarized opinions expressed with free-form text [24]. However, little is known whether this AI predictive capability can expand to more complex ballot formats that involve more options to choose from [5]. Participatory budgeting [25] is one such process put under scrutiny in this article. Here city authorities distribute a public budget by letting citizens propose their own project ideas, which they vote for and often implement themselves [26]. Projects may be pertinent to different impact areas (e.g. environment, culture, welfare), beneficiaries (e.g. elderly, children) and can have different costs [27]. Voters can approve, rank or distribute points over their preferred projects, while winners are elected based on the popularity of the projects (utilitarian greedy) or based on a proportional representation of the voters’ preferences (equal shares or Phragmen’s rule) [28, 29]. Recently though, a reinforcement learning approach has been used to elect winners by assisting voters to reach a consensus [30]. So far, the scope and citizens’ engagement in participatory budgeting campaigns remains to a large extent a one-shot and highly local [31, 11]. What if there was a way to geographically spread such decision-making processes and run them on a regular basis? With such complexity and degree of design freedom, scaling up participatory budgeting turns into the ultimate democratic blueprint to assess capabilities and risks of generative AI voting.
How resilient representative voting outcomes are with generative AI. We hypothesize that generative AI representatives of human voters build up resilience for representative voting outcomes in voting processes of low turnout if these human voters would otherwise abstain or would not have capacity to actively participate (see participation modalities in Figure 1b). In other words, we explore whether the inconsistencies of voting outcomes because of low voting turnout are higher than inconsistencies originated from generative AI representatives of abstained voters. This overall consistency recovery by AI representation defines the resilience of a voting outcome in voting scenarios of low voting turnouts and mixed populations of humans and AI representatives of abstained voters.
Disentangling the role of voting design in generative AI voting. The inconsistencies of generative AI, along with the potential biases explaining these inconsistencies, are systematically studied here for the first time using a novel factorial design based on real-world empirical evidence. It consists of 7 dimensions (see Figure 1b): (i) Real-world voting scenarios - datasets from the 2012, 2016 and 2020 US national elections [32] as well as data from the 2023 participatory budgeting campaign of ‘City Idea’ in Aarau, Switzerland [33]. The latter dataset comes with two voting scenarios, a hypothetical one provided to voters before voting via a survey and the actual voting data. These three scenarios cover a profound range of different number of ballot options and voters to experiment with, see Figure 1b and Section 4.2. (ii) Personal human traits - for each voter, multiple incremental levels of additional information are given as input to large language models to generate ballots. This information includes socio-demographics, political preferences, preferences for project categories and preferences for qualities of voting outcomes (popular vs. fair election of winners, individual vs. collective benefit). These are derived from real-world evidence including survey data collected along the voting campaigns (see Section 4.1, and Table LABEL:table:prologue1 and LABEL:table:prologue2). (iii) Ballot formats - four methods with incremental level of complexity and expressiveness are compared [19, 34]. These include single choice for all voting scenarios, approvals, score (assigning a preference score to each option) and cumulative voting (distributing a number of points over the options) [35, 36, 37] for the participatory budgeting scenarios. (iv) AI models - Three large language models are assessed along with a more mainstream predictive machine learning (ML) model used as a benchmark. GPT 3, GPT 3.5, and Llama3-8B are chosen covering a large spectrum of capabilities in open and closed generative AI [38], see Section 4.1 for more information. (v) Ballot aggregation methods - majority is used for the scenario of US elections. The standard utilitarian greedy, the method of equal shares [39] and the Phragmen’s sequential rule [40] are used for the participatory budgeting scenarios. Utilitarian greedy iterates to elect the next most popular project as long as the budget is not exhausted. It has been the standard method used in real-world. The latter two methods refer to as the fair collective choice based on the principle of proportionality: they elect a combination of projects such that a more proportional representation of voters’ choice is achieved by providing to voters power to decide for an equal part of the available budget [39]. Equal shares is the method that was actually used in City Idea campaign to elect the winners [27], providing a compelling realism for the findings of this study. See section 4.1 for more information. (v) Abstaining models - the voters with low engagement and trust in political processes, as well as low digital literacy represent the ones who are likely to abstain. They are identified using available survey data and grounded to well-established literature [12, 41, 42] (see Section 4.1). (vi) Participation modalities - we assess the consistency of voting scenarios with full and low turnouts of human voters, partial / full AI representation of abstained voters and AI representation of the whole human population. The dimensions of the factorial design are illustrated in Table 1 and the studied combinations are marked with the colored boxes in Figure 1b. This broad spectrum of analysis based on real-world evidence allows us to generalize the findings of the study and make them relevant for a broad spectrum of research communities and policymakers.
@@
Assessing generative AI voting in action. Voting personas are constructed using input prompts of large language models as depicted in Figure 1c. This designed process aims to emulate the three voting scenarios with the different settings of Figure 1b. Each input prompt consists of a standardized description of the voter’s profile (see Section LABEL:sec:study) and an instruction to vote according to the ballot format. The individual and collective real-world choices of humans and AI personas are compared for the first time by measuring the relative number of Condorcet pairwise matches as an indicator of consistency [13, 22, 19]. These consistency values are then becoming the dependent variable to predict using the personal human traits as independent variables (features), fed in a neural network (see Section 4.4). Based on a systematic mapping of human personal traits to cognitive biases as illustrated in Figure 1c (see Section LABEL:sec:bias for more detail), this prediction model causally explains the human traits that contribute to inconsistencies and the potential underlying biases that explain these inconsistencies. This novel analysis is designed to provide a significant conceptual advance in understanding how voting design reinforces or mitigates different AI biases in real-world practice.
@@ 2 Results
1. Fair voting methods to elect winners are more resilient to inconsistencies of AI to accurate estimate human choice, demonstrating a striking underlying win-win relationship: fairer voting outcomes for humans with fairer human representation by AI (Figure 2 and 4). These inconsistencies are particularly prominent in complex ballot formats with a large number of alternatives, while simple majoritarian voting tends to be highly consistent. AI intransitivity across ballot formats is higher than the minimal one of humans, with higher impact though on collective choice over a large number of alternatives (Figure 3).
2. AI representation is more effective over voters who is likely to abstain than an arbitrary voter, in particular under fair collective choice (Figure 4). Abstaining voters result in a representation deficit restored by AI, while AI representation over arbitrary voters mainly has a noise-reduction effect or voting outcome.
3. Features of abstaining voters related to their low engagement, digital literacy and trust explain the consistency of their AI representation and transitivity of ballot formats (Figure 5). Affect and unconscious biases explain the (in)consistency of human-AI choice, while affect and time discounting biases explain transitivity of AI choice over ballot formats.
@@ 2.1 Fair collective choice is resilient to human-AI inconsistencies
Voting design and choice context have an impact on human-AI inconsistencies. Figure 2 illustrates the human-AI consistency in individual and collective choices for single-choice and multi-choice voting. For multi-choice voting, the individual and collective consistency of human and AI choices are measured as the average consistency across various sampled population sizes of 25%, 50%, and 75% (shown separately in Figure LABEL:fig:consistency-biases_actual and LABEL:fig:consistency-biases_survey). The consistency of individual choice remains poor in complex ballot formats with several alternatives. On average, it is 5.2% and 27.15% for the actual and survey voting scenario of City Idea, yet it is 82.3% for the binary majoritarian US elections. GPT 3.5 shows the highest consistency of individual choice among the large language models, which is 4.27% and 4.16% higher than GPT 3 and Llama3-8B. On the contrary, the consistency of collective choice increases by 45.03% in total. Strikingly, the consistency of equal shares and Phragmen’s is on average 68.8%, which is 30.7% higher than utilitarian greedy. Even when reducing the number winners of equal shares to that of utilitarian greedy, the consistency remains 21.4% higher. Compared to large language models, the machine learning model shows 3.3% higher consistency in individual choice and 4.2% higher in collective choice. The consistency values are the maximum ones derived with all personal human traits as shown in Figure 1a. Removing project preferences from the context of AI choice generation results in the highest consistency reduction of 18.1%, whereas political interest leads to the lowest reduction of 3.5%.
@@
Figure 2: Choice by large language models is consistent to humans for single-choice majoritarian elections, however accuracy drops for more complex ballots with larger number of alternatives as in the case of participatory budgeting. Strikingly, accuracy of collective choice is significantly higher than individual choice, particularly for the fairer ballot aggregation rules of equal shares and Phragmén’s. GPT 3.5 shows the highest consistency and Llama3-8B the lowest among the large language models, which though remain inferior to a predictive machine learning model. The mean consistency (y-axis) for different population of voters (10%, 25%, 75% and 100%) in individual and collective choice is shown for different AI models (x-axis), across three real-world voting scenarios: The participatory budgeting campaign of City Idea, (A) actual and (B) survey, as well as (C) the US national elections of 2012, 2016 and 2020. For participatory budgeting, the ballot formats of cumulative/score (left) and approval (right) are shown, including the ballot aggregation methods of equal shares, Phragmén’s and utilitarian greedy. For the actual voting of City Idea, the accuracy of equal shares is calculated for all winners and a controlled number of winners (as many as greedy) for a fairer comparison.
@@
Intransitivity: higher for AI with impact on collective choice among many alternatives. Figure 3 illustrates the transitivity of preferences in different large language models and humans by measuring the consistency of individual and collective choice among different pairs of ballot formats. While human transitivity approaches 97.1% on average, AI transitivity remains at 72.1% for GPT 3.5 and 78.3% for Llama3-8B. In terms of collective choice, equal shares has 12.2% higher consistency than utilitarian greedy. Equals shares has more than 80% consistency among winners from cumulative / score and approval ballots. For the actual voting of City Idea, ‘approval - cumulative’ voting shows the highest transitivity, which is 4.4% higher than ‘single choice - approval’ and 3.2% higher than ‘single choice - cumulative’. However, for the survey of City Idea, ‘single choice - cumulative’ shows the highest transitivity, which is 25.2% higher than ‘single choice - approval’ and 15.1% higher than ‘approval - cumulative’.
Figure 3: Intransitivity of AI across different pairs of ballot formats is higher than the one of humans, which remains negligible. AI intransitivities have a higher influence on the consistency of voting outcomes over a large number of alternatives. Llama3-8B predicts ballots that are not very diverse and selects a limited set of projects, which results in higher transitivity compared to other language models. Equal shares and Phragmén’s also show here higher capacity to mitigate the ballot intransitivities. It achieves more than 80% consistency between cumulative and approval ballots. The consistency (y-axis) in individual choice among different pairs of ballot formats (x-axis) is shown for different large language models, humans and the two voting scenarios in the participatory budgeting campaign of City Idea: (A) actual vs. (B) survey. Mean consistency values are calculated across randomly sampled population of 25%, 50%, and 75%.